OpenStudy (anonymous):

Please Help Multiply (x – 4)(x2 + 5x + 6). x3 + 7x2 – 20x – 24 x3 + x2 – 14x – 24 x3 + 9x2 – 26x – 24 x3 + 5x2 + 10x – 24

OpenStudy (anonymous):

∫ [(x - 4) /(x² - 2x - 3x + 6)] dx = ∫ {(x - 4) /[x(x - 2) - 3(x - 2)]} dx = ∫ {(x - 4) /[(x - 2)(x - 3)]} dx = then decompose the integrand into partial fractions: (x - 4) /[(x - 2)(x - 3)] = A/(x - 2) + B/(x - 3) (x - 4) /[(x - 2)(x - 3)] = [A(x - 3) + B(x - 2)] /[(x - 2)(x - 3)] x - 4 = A(x - 3) + B(x - 2) x - 4 = Ax - 3A + Bx - 2B x - 4 = (A + B)x + (- 3A - 2B) A + B = 1 - 3A - 2B = - 4 A = 1 - B 3A + 2B = 4 A = 1 - B 3(1 - B) + 2B = 4 A = 1 - B 3 - 3B + 2B = 4 A = 1 - B - B = 4 - 3 A = 1 - B - B = 1 A = 1 - (- 1) = 1 + 1 = 2 B = - 1 hence: (x - 4) /[(x - 2)(x - 3)] = A/(x - 2) + B/(x - 3) = 2/(x - 2) - 1/(x - 3) thus the integral becomes: ∫ [(x - 4) /(x² - 5x + 6)] dx = ∫ {[2/(x - 2)] - [1/(x - 3)]} dx = break it up pulling constants out: 2 ∫ [1 /(x - 2)] dx - ∫ [1 /(x - 3)] dx = 2 ln |x - 2| - ln |x - 3| + C (antiderivative) having the antiderivative, plug in the bounds: 1 ∫ [(x - 4) /(x² - 5x + 6)] dx = (2 ln |1 - 2| - ln |1 - 3|) - (2 ln |0 - 2| - ln |0 - 3|) = 0 2 ln |- 1| - ln |- 2| - 2 ln |- 2| + ln |- 3| = 2 ln (1) - ln (2) - 2 ln (2) + ln (3) = 2(0) - 3 ln (2) + ln (3) = (recalling logarithm properties) ln (3) - ln (2³) = ln (3) - ln (8) = ln (3/8) ≈ - 0.980829

OpenStudy (anonymous):

I don't exactly understand that

OpenStudy (minimato):